OpenStudy (anonymous):

Consider the following data points (1, 4) (2, 9) (3, 6) (4, 10) Use partial derivatives to obtain the formula for the best least-squares fit to the data points.

OpenStudy (anonymous):

It will look like a square

OpenStudy (anonymous):

i got y(x)=-30.5x+99.5 which is wrong..

myininaya (myininaya):

so far i got \[L(x)=\frac{5}{3}(x-1)(x-2)(x-3)- \] \[3(x-1)(x-2)(x-4)-\frac{9}{2}(x-1)(x-3)(x-4)-\frac{2}{3}(x-2)(x-3)(x-4) \]

myininaya (myininaya):

this is my work i hope you can read some it lol

myininaya (myininaya):

i did it with Lagrange polynomials